OpenStudy (anonymous):

please help, i don't understand this at all. an explanation would be really appreciated! A surveyor measures the angle of elevation at 20° for a point 5 miles away. What is the vertical change in elevation from the point where the surveyor is standing to the point 5 miles away?

OpenStudy (stamp):



OpenStudy (anonymous):

thanks!! how does that relate to the problem can you explain that part? sorry

OpenStudy (stamp):

You are currently doing right triangle trigonometry in class, you have a right triangle. Use your trigonometry skills to solve for y.

OpenStudy (anonymous):

how can i find y from here?

OpenStudy (anonymous):

@stamp

OpenStudy (anonymous):

From the triangle, you get the following ratio: \[\tan20^0=\frac{y}{5} \]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

so would that be the answer?

OpenStudy (anonymous):

@sithsandgiggles

OpenStudy (stamp):

@erin512 You need to be slightly independent when it comes to doing your homework. Read your textbook, I am serious, read the whole section on this topic, and then come back if you still do not understand.

OpenStudy (anonymous):

ok, i dont understand all the stuff in the textbook, i dont want u to give me the answer i just need help on how to find it. the textbook doesn't explain it step by step or even work the problem

OpenStudy (anonymous):

\[\tan20^0=\frac{y}{5}\\ y=5\tan20^0\] All you have to do is use a calculator, unless you have tan(20°) memorized.

OpenStudy (stamp):

It is a bit difficult to help you any further without giving you the answer. I am guessing you are in middle school taking an Algebra I class? Or is this Geometry.

OpenStudy (anonymous):

0.36 is tan(20degrees) right?

OpenStudy (anonymous):

this is trig

OpenStudy (anonymous):

oh or is it tan20^0

OpenStudy (anonymous):

cause thats 1.56

OpenStudy (stamp):

@erin512 Ok so I assume they do not place you in trigonometry without a background in algebra. Once you have your ratio, you solve for the variable y (see @SithsAndGiggles reply). \[y=5tan(20°)\]

OpenStudy (stamp):

It is 20°, Sith was using a 0 in lieu of a degree symbol.

OpenStudy (anonymous):

i got 1.82

OpenStudy (anonymous):

is that right?

OpenStudy (stamp):

@erin512 I do not know, I often do not use calculators and I do not have one on my person.

OpenStudy (anonymous):

http://mathway.com

OpenStudy (stamp):

@erin512 It is 1.82, you are correct.

OpenStudy (anonymous):

thank you :) is there another step or is that it?

OpenStudy (stamp):

Make sure to include your unit, miles. The answer is 1.82 miles. See the updated picture to visualize what we just accomplished. 

OpenStudy (anonymous):

thank you so much, i understand it a bit better now

OpenStudy (anonymous):

so y=1.82 miles basically

OpenStudy (stamp):

@erin512 What is the vertical change in elevation from the point where the surveyor is standing to the point 5 miles away? The vertical change in elevation y is 1.82 miles.

OpenStudy (anonymous):

thanks so much!!!